OpenStudy (anonymous):

Billye ran to town at 8 kilometers per hour and then walked back home at 3 kilometers per hour. How far was it to town if the round trip took 11 hours?

OpenStudy (anonymous):

got anything?

OpenStudy (mysesshou):

sorry computer slow. many windows open, jumped to another screen while loaded. -_- ok. give sec

OpenStudy (anonymous):

K

OpenStudy (mysesshou):

8km/h to town 3km/h to home distance = ? time = 11h distance = rate*time Let D = the distance to town D/(8km/h) + D/(3km/h) = time (3km/h)* D + (8km/h) *D = time *(8 * 3) 11 km/h * D = 24km/h * 11hr D = 24 km to the town to check: time = distance/rate 24km/(8km/h) + 24km/(3km/h) = 3h + 8h = 11h how does that seem?

OpenStudy (mysesshou):

ok. have to go for now. have to get home before the thunderstorm rolls in fully. :) good luck with your problems.

OpenStudy (anonymous):

yup